OpenStudy (anonymous):

if someone is 70 inches tall and need a 210 centimeter skis what would a person that 63 inches need

OpenStudy (texaschic101):

70/210 = 63/x (70 in to 210 cm = 63 in to x cm) now cross multiply (70)(x) = (63)(210) 70x = 13230 -- now divide by 70 x = 13230/70 x = 189 63 inch person needs 189 cm skis